### v1

# Progressive-Proximity Bit-Flipping for Decoding Surface Codes

Abstract
Topological quantum codes, such as toric and surface codes, are excellent candidates for hardware implementation due to their robustness against errors and their local interactions between qubits. However, decoding these codes efficiently remains a challenge: existing decoders often fall short of meeting requirements such as having low computational complexity (ideally linear in the code’s blocklength), low decoding latency, and low power consumption. In this paper we propose a novel bit-flipping (BF) decoder tailored for toric and surface codes. We introduce the proximity vector as a heuristic metric for flipping bits, and we develop a new subroutine for correcting a particular class of harmful degenerate errors. Our algorithm achieves linear complexity growth and it can be efficiently implemented as it only involves simple operations such as bit-wise additions, quasi-cyclic permutations and vector-matrix multiplications. The proposed decoder shows a decoding threshold of $7.5\%$ for the 2D toric code and $7\%$ for the rotated planar code over the binary symmetric channel.
Index Terms:
Surface codes, topological codes, bit flipping, decoding algorithm, quantum error correction.
I Introduction
Quantum computers make use of the principles of quantum mechanics to perform computations. Quantum states are fragile and very sensitive to errors, and thus it is crucial to implement quantum error correction techniques to protect quantum information. A very important class of quantum codes are topological codes, specifically surface and toric codes [1], as they can be implemented on a planar quantum chip. Decoding of surface codes is typically performed using the minimum-weight-perfect-matching (MWPM) decoder [2]; although MWPM provides excellent decoding performance, its computational complexity makes it infeasible for large-scale implementation. Because of this, many alternative approaches have been studied to lower the decoding complexity of surface codes. The most promising alternative is the union-find (UF) decoder [3], although many other classes of decoders have been proposed, such as tensor network decoders [4] (which are implementations of maximum likelihood decoding), re-normalization group decoders [5], neural network based decoders [6], maxSAT decoders [7] and cellular-automaton decoders [8]. On the other hand, message passing decoders such as belief propagation (BP) have shown to be effective for decoding surface codes when paired with MWPM and UF [9], or with post-processing techniques such as ordered statistics decoding (BP-OSD) [10]. Also, serial scheduling and re-initialization of BP has been shown to provide good decoding performance [11].
Bit flipping (BF) decoders are a class of iterative decoding algorithms known to be very fast and efficient [12], although generally they provide lower performance than BP. In general, BF decoders do not perform well on surface codes, mainly because of the very low column weight of the parity check matrix, and also because of error degeneracy, which refers to the property of quantum codes where multiple error patterns of the same weight can correspond to the same syndrome. Nevertheless, because of its extremely low complexity, BF is still attractive in the scenario of decoding topological codes: indeed, the latency constraint that a quantum decoder should meet is very tight, to prevent the so-called backlog problem [13]. Moreover, it is essential that the decoder has a low power consumption, as it has to be embedded in a cryogenic environment with a strict power budget [14].
In this paper, we develop a BF algorithm which is capable of decoding surface codes. In the proposed approach, instead of considering the number of unsatisfied checks as in conventional BF, each bit is assigned a heuristic weight which is the entry of what we call proximity vector. The proximity vector is the sum of different contributions called individual influences, and each individual influence is associated with an unsatisfied check. Ultimately, bits connected with checks with low entries in the proximity vector will be flipped first, while bits connected with checks with high entries in the proximity vector will be flipped last. After each flip, the proximity vector is updated in an efficient manner. To deal with high-weight consecutive errors, i.e., errors occurring on a set of adjacent qubits, which are particularly harmful for iterative decoders, we design an iterative matching procedure that runs after BF and that is able to correct these errors. We show that our decoder has an asymptotic complexity of $\mathcal{O}(n)$, where $n$ is the code’s blocklength, and we provide simulation results that show a comparison, in terms of performance, with MWPM, UF and traditional BF, as well as the average number of decoding iterations as a function of the blocklength.
The rest of the paper is organized as follows. In Section II we introduce the preliminaries of quantum error correction. In Section III we present an overview of the main decoding algorithms for topological codes. In Section IV we define the proximity vector and how it can be computed efficiently. Section V provides a detailed description of our proposed decoder. In Section VI we carry out a complexity analysis and a comparison with other state-of-art decoders. Finally, Section VII presents simulation results.
II Preliminaries
II-A Stabilizer formalism
Consider the $n$-fold Pauli group,
$\mathcal{G}_{n}\triangleq\{cB_{1}\otimes\cdots\otimes cB_{n}:c\in\{\pm 1,\pm i\},B_{j}\in\{I,X,Y,Z\}\},$
where $I=\begin{bmatrix}1&0\\ 0&1\end{bmatrix},X=\begin{bmatrix}0&1\\ 1&0\end{bmatrix},Y=\begin{bmatrix}0&-i\\ i&0\end{bmatrix},Z=\begin{bmatrix}1&0\\ 0&-1\end{bmatrix}$. Every non-identity Pauli operator $P\in\mathcal{G}_{n}$ has eigenvalues ±$\pm 1$ and any two Pauli operators in $\mathcal{G}_{n}$ either commute or anti-commute with each other. The weight of a Pauli operator $P$ is defined as the number of non-identity elements in the tensor product.
By dropping the phase factor $c$, the Pauli group $\mathcal{G}_{n}$ is isomorphic to $\mathbb{F}_{2}^{2n}$ [10], such that
[…]
In this representation, the commutation relation between two Pauli operators $\mathbf{p}_{1}=(\mathbf{x_{1}}|\mathbf{z_{1}})$ and $\mathbf{p}_{2}=(\mathbf{x_{2}}|\mathbf{z_{2}})$ can be computed by checking that the symplectic inner product is zero:
$\mathbf{x_{1}}\mathbf{z_{2}}^{T}+\mathbf{z_{1}}\mathbf{x_{2}}^{T}=0\ \mod 2.$ (2)
A stabilizer group $\mathcal{S}$ is an Abelian subgroup of $\mathcal{G}_{n}$, and an $\llbracket n,k,d\rrbracket$ stabilizer code is a $2^{k}$-dimensional subspace $\mathcal{C}$ of the Hilbert space $(\mathbb{C}^{2})^{\otimes n}$ that satisfies the condition $S_{i}\ket{\Psi}=\ket{\Psi},\ \forall\ S_{i}\in\mathcal{S},\ket{\Psi}\in\mathcal{C}$. Thus, the code $\mathcal{C}$ is defined as the common $+1$-eigenspace of the stabilizer group $\mathcal{S}$. The stabilizer group $\mathcal{S}$ is generated by a set of $n-k$ independent generators $S_{1},...,S_{n-k}$ that can be represented using a matrix $\mathbf{S}$, called the stabilizer matrix, whose $(i,j)$ element is given by the Pauli operator corresponding to the $j$-th qubit in the $i$-th stabilizer. The minimum distance $d$ is defined as the minimum weight of an element of $\mathcal{S}$.
Applying the mapping (1) to the stabilizer matrix $\mathbf{S}$, we obtain a $(n-k)\times 2n$ binary matrix:
$\mathbf{H}=[\mathbf{H}_{X}\ |\ \mathbf{H}_{Z}]$ (3)
which we call the parity check matrix of $\mathcal{C}$. Since the corresponding stabilizers of $\mathbf{S}$ commute with each other, it is easy to check using (2) that
$\mathbf{H}_{X}\mathbf{H}_{Z}^{T}+\mathbf{H}_{Z}\mathbf{H}_{X}^{T}=\mathbf{0}\ \mod 2.$ (4)
II-B Error model and syndrome
In the depolarizing error model, errors are Pauli operators such that $\mathbf{e}\in\{I,X,Z,Y\}^{\otimes n}$; each error on the $i$-th qubit $e_{i}$ can either be a bit flip ($X$), a phase flip ($Z$) or both ($Y$), each with probability $\epsilon/3$, while the probability of no error ($I$) is equal to $1-\epsilon$. Using the Pauli-to-binary mapping of (1), a Pauli error $\mathbf{e}$ can be also be modeled as a $1\times 2n$ binary vector $\mathbf{e}=[\mathbf{e}_{X}\ \mathbf{e}_{Z}]$. Given a Pauli error $\mathbf{e}$, the corresponding syndrome $\mathbf{s}\in\{0,1\}^{n-k}$ can be computed using the symplectic inner product such that
$\mathbf{s}=\mathbf{e}_{X}\mathbf{H}_{Z}^{T}+\mathbf{e}_{Z}\mathbf{H}_{X}^{T}\ \mod 2.$ (5)
Obviously, if $\mathbf{e}\in\mathcal{S}$ we have $\mathbf{s}=\mathbf{0}$. More generally, any Pauli error can be expressed as a combination of a true error and a stabilizer such that $\mathbf{e}=\mathbf{e}_{t}+\mathbf{S}_{i}$, with $\mathbf{S}_{i}\in\mathcal{S}$; since, by definition, elements of the stabilizer group $\mathcal{S}$ commute with each other (thus, their syndrome is zero), the syndrome $\mathbf{s}$ is only dependent on the true error $\mathbf{e}_{t}$; nonetheless, this also means that for every true error $\mathbf{e}_{t}$, there exist a set of Pauli errors $\mathcal{E}=\mathbf{e}_{t}+\mathcal{S}$ that will lead to the same syndrome. This phenomenon is known as error degeneracy. It is well known that, if the minimum distance of a stabilizer code is much higher than the weight of its stabilizer elements, there will be many degenerate errors of the same weight [15].
II-C Calderbank-Shor-Steane codes
An $\llbracket n,k_{X}-k_{Z},d\rrbracket$ Calderbank-Shor-Steane (CSS) code $\mathcal{C}$ is a stabilizer code constructed using two classical $[n,k_{X},d_{X}]$ and $[n,k_{Z},d_{Z}]$ codes $C_{X}$ and $C_{Z}$, respectively, where $d\geq\leavevmode\nobreak\ \mathrm{min}\{d_{X},d_{Z}\}$ and $C_{Z}\subset C_{X}$ [16]. Note that $k_{X}$, $k_{Z}$ and $d_{X}$, $d_{Z}$ correspond to the dimensions and minimum distances of $C_{Z}$ and $C_{X}$, respectively. The stabilizer matrix of the CSS code $\mathcal{C}$ has the form
$\mathbf{H}=\begin{bmatrix}\mathbf{H}_{Z}&\mathbf{0}\\ \mathbf{0}&\mathbf{H}_{X}\end{bmatrix},$ (6)
where $\mathbf{H}_{X}$ and $\mathbf{H}_{Z}$ are the parity check matrices of $C_{X}$ and $C_{Z}$, respectively, and the commutativity condition of (4) reduces to $\mathbf{H}_{X}\mathbf{H}_{Z}^{T}=\mathbf{0}\mod 2$. To correct depolarizing errors on the qubits, a syndrome $\mathbf{s}$ is computed such that
$\mathbf{s}=[\mathbf{s}_{X}\ \mathbf{s}_{Z}],$ (7)
where $\mathbf{s}_{X}=\mathbf{e}_{X}\mathbf{H}_{Z}^{T}\mod 2$ and $\mathbf{s}_{Z}=\mathbf{e}_{Z}\mathbf{H}_{X}^{T}\mod 2$. Because of the structure of $\mathbf{H}$, $X$ and $Z$ errors can be corrected independently using $\mathbf{H}_{Z}$ and $\mathbf{H}_{X}$, respectively. In this paper we consider a bit-flip channel, where each qubit experiences an $X$ error with probability $p$, and remains correct with probability $1-p$. In other words, we fix $\mathbf{e}_{Z}=\mathbf{0}$, while elements in $\mathbf{e}_{X}$ can be 1 with probability $p$ or 0 with probability $1-p$. Hence we only consider $\mathbf{H}_{Z}$ and $\mathbf{s}_{X}$, which for simplicity we will denote as $\mathbf{H}$ and $\mathbf{s}$, respectively. Note that the bit-flip channel that we consider and the depolarizing error model are closely related: indeed, assuming that $X$ and $Z$ errors are decoded separately, we can model the depolarizing channel as two binary symmetric channels with probability of error $p_{x}=p_{z}=\frac{2}{3}\epsilon$; assuming that $d_{X}=d_{Z}$, is possible to switch from the logical error rate curve over a binary symmetric channel to the logical error rate under depolarizing noise by re-scaling it of a factor of $3/2$ [17].
It is convenient to introduce the notion of Tanner graph [18]. A Tanner graph is a bipartite graph defined from $\mathbf{H}$, such that it has two sets of nodes $V=\{v_{1},v_{2},...,v_{n}\}$ and $C=\{c_{1},c_{2},...,c_{n-k}\}$ called variable nodes and check nodes, respectively, and there is an edge between $v_{j}$ and $c_{i}$ if and only if $h_{i,j}=1$. A check node $c_{i}$ is said to be satisfied if $s_{i}=0$, and unsatisfied if $s_{i}=1$. The degree of a variable or check node is defined as the number of its incident edges. We define the distance between two nodes $i$ and $j$ to be the number of variable nodes belonging to the shortest path between $i$ and $j$.
II-D Toric and surface codes
Toric codes [1] are a widely known class of CSS codes. These are derived from the tessellation of the topological surface of a torus in squares, in such a way as to form a lattice. Generally, a toric code is characterized by a $L\times L$ lattice, where $L$ is the size of the horizontal (or vertical) dimension. To an $L\times L$ lattice corresponds a $\llbracket 2L^{2},2,L\rrbracket$ quantum CSS code. The $\mathbf{H}_{X}$ matrix is the incidence matrix between vertices (check nodes) and edges (variable nodes), and the $\mathbf{H}_{Z}$ matrix is the incidence matrix between squares (check nodes) and edges (variable nodes). The two Tanner graphs corresponding to $\mathbf{H}_{X}$ and $\mathbf{H}_{Z}$ are identical and they obviously correspond to a torus as well, as depicted in Fig. 1.
Surface codes are similar to toric codes, except that the lattice has now open boundaries; in such a way, the code can be implemented on a two-dimensional quantum chip [19]. Although many variants of surface codes have been studied, we focus on the $\llbracket L^{2},1,L\rrbracket$ rotated planar codes [19], which are obtained by a $45\degree$ rotation of an $L\times L$ lattice, after the pruning of unnecessary qubits.
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
Figure 1: Tanner graph of the $\llbracket 18,2,3\rrbracket$ toric code. Black values correspond to variable node labels, while red values are check node labels.
III Decoding of toric and surface codes
In this section we summarize the state-of-the-art decoders for toric and surface codes, and motivate our contribution.
III-A Minimum weight perfect matching
The decoding of toric and surface codes is usually carried out using the MWPM algorithm [2]. After measuring the syndrome, a complete graph is constructed, where each vertex represents an unsatisfied check; each edge is assigned a weight, which is the minimum number of qubits separating the two checks connected by the edge; this weight assignment is done using Dijkstra’s algorithm [20]. Then, the blossom algorithm [21] is run on this graph, with the goal of finding the minimum weight perfect matching. A matching on a graph is a subset of edges such that any two edges in this subset do not share any common vertex; a perfect matching is a matching such that every vertex in the graph is connected to exactly one edge. A minimum weight perfect matching is a perfect matching such that the sum of all the weights of the edges of the matching is minimum among all the possible perfect matchings. Since the overall complexity of the MWPM decoder on an $L\times L$ lattice is $\mathcal{O}(L^{6}\log L)$ [22], many efficient implementations have been proposed to reduce this complexity: for instance, in [2] a fully parallel implementation is proposed, able to run with a complexity of $\mathcal{O}(1)$, given a uniform 2-D array of finite speed processing elements and an external memory able to store all the detection events and matching data for the entire duration of a hypothetical computation, while Higgott et al. proposed, in [22], that each unsatisfied check could be restricted to be matched within a local neighborhood, or that Dijkstra’s algorithm could be optimized to avoid an all-to-all search [23].
III-B Union-find
The main competitor of MWPM is the UF decoder [3]. This decoding scheme is also divided into two parts: first, the syndrome validation process is carried out, where clusters are defined to initially correspond to each unsatisfied check, and are then iteratively grown to absorb the nearest check neighbors. A cluster stops growing when it contains an even number of unsatisfied checks, and two clusters merge together if they touch each other. After all the clusters have stopped growing, the peeling decoder is run on each cluster; it is possible to treat errors as erasures because the error location is known after the syndrome validation step, given that each cluster is a spanning tree; otherwise, clusters are split in order to obtain spanning trees [24]. The complexity of the UF decoder, in its efficient implementation proposed in [3], is almost linear in the blocklength.
III-C Message-passing decoders
Message passing decoders such as BP are, in principle, not suitable for topological codes, mainly because of the presence of weight-4 symmetric stabilizers that behave as trapping sets for the decoder [25]; because of this, several techniques have been developed to improve the performance of BP on these codes. One of the most common approaches is to apply the well-known ordered statistics technique after several rounds of BP [10]; this decoder is known as BP-OSD and has a complexity of $\mathcal{O}(n^{3})$, thus making it infeasible for large-scale implementation. More recent contributions have shown that a serial scheduling of the messages, along with modifications of the update rules and re-initialization, lead to a significant increase of the performance at the cost of increasing the decoding latency [11, 26].
III-D Motivation for our work
For the majority of decoders proposed for topological codes, low decoding error probability comes with the price of increased decoding complexity. MWPM, for instance, has a complexity that is cubic in the blocklength and it is known to be infeasible in practice; the efficient implementations available, in particular those of [22] and [23], offer a significant speedup that still vanishes for large blocklength. The UF decoder, on the other hand, is able to offer an almost-linear time complexity by exploiting an efficient implementation, and its performance is somewhat close to that of MWPM; nevertheless, parallel implementations of this decoder come with several challenges: in [27], for instance, it is shown that the worst-case complexity rises to $\mathcal{O}(L^{3})$ when the decoder is implemented using parallel computational units that exchange messages with their nearest neighbors. Indeed, besides low asymptotic complexity, it is important that a decoder also has low decoding latency. In this sense, message-passing decoders such as BP are known to be excellent, as the update rules are local and equal for each node type, and can be easily implemented in parallel. Nevertheless, algorithms like BP-OSD fail to achieve linear complexity.
In this scenario, it seems reasonable to revisit the BF decoder, as it not only exhibits a complexity linear in $n$, but also utilizes extremely simple update rules that make it much faster than BP. However, because of the low degree of variable nodes in the Tanner graph of the considered quantum codes (toric codes are regular with variable degree of 2, while surface codes also have degree-1 nodes), and because of error degeneracy, we need to make significant modifications to the traditional BF algorithm in order to decode toric and surface codes. We will show that these modifications do not impact on the asymptotic complexity of the decoder, and that they significantly improve the performance of the decoder comparing to the one of the traditional BF; we also compare the performance of our algorithm with the ones of MWPM and UF.
IV The proximity vector
In this section we introduce the proximity vector, a vector that the decoder uses as a bit flipping criterion. We describe the rationale behind it, and we describe how it can be efficiently computed while decoding.
The proximity vector is a heuristic weight assignment to the variable nodes and check nodes in the Tanner graph, and we define it to be the sum of multiple contributions which we call proximity influences, which we describe hereafter. Let $c_{j}$ be an unsatisfied check; we want variable and check nodes neighboring $c_{j}$ to have the highest weights, while the weights should decrease with increasing distance from $c_{j}$. A simple way to achieve this is by computing the proximity influence recursively, as follows:
Definition 1.
Let $c_{j}$ be an unsatisfied check, and let all the other check nodes be satisfied. We define $\boldsymbol{\gamma}^{(0)}(c_{j})$ to be a $1\times m$ vector such that
$\gamma_{i}^{(0)}(c_{j})=\begin{cases}1\ \mathrm{for}\ i=j\\ 0\ \mathrm{otherwise}\ ,\end{cases}$ (8)
and let $\boldsymbol{\nu}^{(0)}(c_{j})$ be a $1\times n$ vector such that
$\boldsymbol{\nu}^{(0)}=\boldsymbol{\gamma}^{(0)}\cdot\mathbf{H}\ .$ (9)
Then, let $\boldsymbol{\gamma}^{(\ell)}(c_{j})$ and $\boldsymbol{\nu}^{(\ell)}(c_{j})$ be defined recursively as
$\begin{cases}\boldsymbol{\gamma}^{(\ell)}(c_{j})=\boldsymbol{\nu}^{(\ell-1)}(c_{j})\cdot\mathbf{H}^{T}\\ \boldsymbol{\nu}^{(\ell)}(c_{j})=\boldsymbol{\gamma}^{(\ell)}(c_{j})\cdot\mathbf{H}\ ,\end{cases}$ (10)
for $\ell=1,2,...,D$. We call $\boldsymbol{\nu}^{(\ell)}(c_{j})$ the proximity influence of $c_{j}$ on qubits (or variable nodes) of depth $\ell$, and $\boldsymbol{\gamma}^{(\ell)}(c_{j})$ the proximity influence of $c_{j}$ on checks of depth $\ell$.
We then combine the proximity influence of each unsatisfied check node to obtain the proximity vectors by adding, for each variable and check node, the values of each proximity influence.
Definition 2.
Let $c_{1},...,c_{m}$ be all the unsatisfied check nodes, and let $\boldsymbol{\nu}^{(D)}(c_{1}),...,\boldsymbol{\nu}^{(D)}(c_{m})$ and $\boldsymbol{\gamma}^{(D)}(c_{1}),...,\boldsymbol{\gamma}^{(D)}(c_{m})$ be the respective proximity influences on qubits and checks. We define $\boldsymbol{\nu}^{(D)}$ and $\boldsymbol{\gamma}^{(D)}$ to be the proximity vectors on qubits and check nodes of depth $D$, respectively, such that:
$\begin{cases}\boldsymbol{\nu}^{(D)}=\sum_{i=1}^{m}\boldsymbol{\nu}^{(D)}(c_{i})\\ \boldsymbol{\gamma}^{(D)}=\sum_{i=1}^{m}\boldsymbol{\gamma}^{(D)}(c_{i})\end{cases}.$ (11)
It is also possible to define $\boldsymbol{\nu}^{(D)}$ and $\boldsymbol{\gamma}^{(D)}$ by setting $\boldsymbol{\gamma}^{(0)}=\mathbf{s}$, and then using (9) and (10) directly.
Since the value $D$ will be fixed in the rest of the paper, we simply refer to the influences as $\boldsymbol{\nu}(c_{j})$ and $\boldsymbol{\gamma}(c_{j})$, and to the vectors as $\boldsymbol{\nu}$ and $\boldsymbol{\gamma}$. By construction, the values of the influences $\boldsymbol{\nu}(c_{j})$ and $\boldsymbol{\gamma}(c_{j})$ are highest for the nearest neighbors of $c_{j}$, and decrease with increasing distance of a variable (check) node from $c_{j}$. The proximity vector on qubits $\boldsymbol{\nu}$ and on checks $\boldsymbol{\gamma}$ is the natural superposition of all of the influences exerted by all of the unsatisfied checks. As a result, variable and check nodes which are more distant from unsatisfied nodes (we say more isolated) will have a lower weight assigned, while variable and check nodes that are near to many unsatisfied checks will have higher weights assigned. Note that this definition of proximity vector is heuristic, and alternative definitions can be provided to capture the individual influence of a check node. Our definition of proximity vector is convenient for the reason that it only involves integer values, thus it can be stored using a low fixed (and low) number of bits.
IV-1 Efficient computation of the proximity vector
In our decoder, it is essential to update the proximity vectors on checks and qubits after each flip: in other words, after a bit is flipped, its neighboring checks will be satisfied [^1] Because of the nature of our decoder, each bit flip never generates new unsatisfied checks., and their influence should be removed from the proximity vector. Specifically, if the flipping of variable node $v_{i}$ causes its two neighboring check nodes $c_{j}$ and $c_{k}$ to become satisfied, the updated proximity vectors shall be
$\begin{cases}\boldsymbol{\gamma}^{\prime}=\boldsymbol{\gamma}-(\boldsymbol{\gamma}(c_{j})+\boldsymbol{\gamma}(c_{k}))\\ \boldsymbol{\nu}^{\prime}=\boldsymbol{\nu}-(\boldsymbol{\nu}(c_{j})+\boldsymbol{\nu}(c_{k})).\\ \end{cases}$ (12)
To take into account this update of the vector after each flip, one could in principle recompute (11), setting $\boldsymbol{\gamma}^{(0)}=\mathbf{s}^{\prime}$, where $\mathbf{s}^{\prime}$ is the updated syndrome after the bit flip; however, this would negatively impact the complexity of the algorithm, as each bit flip would require an additional $\ell(nd_{v}+md_{c})$ operations (multiplication of sparse matrices).
Instead, we exploit the quasi-cyclic property of the toric code to make the procedure more efficient. The idea is to pre-compute the proximity influence of an arbitrary check node, say $\boldsymbol{\gamma}(c_{1}),\boldsymbol{\nu}(c_{1})$, store it prior to the decoding process, and then compute online $\boldsymbol{\gamma}(c_{i}),\boldsymbol{\nu}(c_{i})$ when needed, by appropriately permuting $\boldsymbol{\gamma}(c_{1}),\boldsymbol{\nu}(c_{1})$. Assume that we have an $L\times L$ toric code. We can label the variable nodes with non-negative integers row-wise in increasing order (so that the first row is labeled from 00 to $L-1$, the second row is labeled from $L$ to $12L-$ 1, and so on), and we can do the same for the check nodes. An example is provided in Fig. 1 for the case $L=3$. There will be a total of $2L^{2}$ variable nodes and $L^{2}$ check nodes. Assume that we have already stored $\boldsymbol{\nu}(c_{i})$ and $\boldsymbol{\gamma}(c_{i})$ for some arbitrary check node $c_{i}$, and that we now wish to compute $\boldsymbol{\nu}(c_{j})$ and $\boldsymbol{\gamma}(c_{j})$, for some $j\neq i$, by appropriately permuting $\boldsymbol{\nu}(c_{i})$ and $\boldsymbol{\gamma}(c_{i})$. By exploiting the indexing we have defined for the check and variable nodes, we can define a vertical shift $\sigma_{y}$ and an horizontal shift $\sigma_{x}$ in the following manner:
$\begin{cases}\sigma_{y}=\lfloor(j-i)/L\rfloor\\ \sigma_{x}=j-i\mod L\ \ ,\end{cases}$ (13)
assuming that $i<j$. We can express the index transformation in a closed form, using $\sigma_{x}$ and $\sigma_{y}$. Let $k_{c}\in[1,L^{2}]$ and $k_{v}\in[1,2L^{2}]$ be the indices of the check and variable nodes, respectively. We can express a coordinate shift using two linear maps from $k_{c}$ to $k^{\prime}_{c}$ and from $k_{v}$ to $k^{\prime}_{v}$ respectively, such that
$\begin{cases}k^{\prime}_{c}=\bigl{\{}\{k_{c}+\sigma_{x}\}\mod L\ +\\ \hskip 25.0ptL\lfloor(k_{c}-1)/L\rfloor+\sigma_{y}L\bigr{\}}\mod L^{2}\\ k^{\prime}_{v}=\bigl{\{}\{k_{v}+\sigma_{x}\}\mod L\ +\\ \hskip 25.0ptL\lfloor(k_{v}-1)/L\rfloor+2\sigma_{y}L\bigr{\}}\mod 2L^{2}\\ \end{cases}$ (14)
Hence, we have $\gamma_{k^{\prime}_{c}}^{\ell}(c_{j})=\gamma_{k_{c}}^{\ell}(c_{i})$ and $\nu_{k^{\prime}_{c}}^{\ell}(c_{j})=\nu_{k_{c}}^{\ell}(c_{i})$. An example of this transformation is illustrated in Fig. 2. In Fig 1(a), the proximity influence of $c_{4}$ of depth 1, namely $\boldsymbol{\nu}^{\ell}(c_{4})$ and $\boldsymbol{\gamma}^{\ell}(c_{4})$, is assumed to have been pre-computed. In Fig. 1(b), the influence is shifted to compute $\boldsymbol{\nu}^{\ell}(c_{8})$ and $\boldsymbol{\gamma}^{\ell}(c_{8})$; in the example, $\sigma_{x}=\sigma_{y}=1$. The reader can verify that, using (14), the variable node $4$ is mapped to $11$, variable node $6$ to $13$ and so on; similarly, the influence on the checks is shifted. We utilize the $\mathtt{Shift\_influence}$ subroutine to apply the coordinate permutation of (14); the subroutine, described in Algorithm 1, takes as input the index of a check node $c_{j}$, and applies (14) to $\boldsymbol{\gamma}(c_{1})$ and $\boldsymbol{\nu}(c_{1})$ to obtain $\boldsymbol{\gamma}(c_{j})$ and $\boldsymbol{\nu}(c_{j})$.
Algorithm 1 $\mathtt{Shift\_influence}$
Input: $j$
Global: $\boldsymbol{\gamma}(c_{1})$, $\boldsymbol{\nu}(c_{1})$, $L$
Output: $\boldsymbol{\gamma}(c_{j})$, $\boldsymbol{\nu}(c_{j})$
1:$\mathbf{k}_{v}\leftarrow[1,2L^{2}]$
$\mathbf{k}_{c}\leftarrow[1,L^{2}]$
3:Compute $\sigma_{x},\sigma_{y}$ using (13)
4:Compute $\mathbf{k}^{\prime}_{v},\mathbf{k}^{\prime}_{c}$ using (14)
5:$\boldsymbol{\nu}_{\mathbf{k}^{\prime}_{v}}(c_{j})\leftarrow\boldsymbol{\nu}_{\mathbf{k}_{v}}(c_{1})$
6:$\boldsymbol{\gamma}_{\mathbf{k}^{\prime}_{c}}(c_{j})\leftarrow\boldsymbol{\gamma}_{\mathbf{k}_{c}}(c_{1})$
7:return
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
(a) The depth-1 influence of $c_{4}$ (highlighted in red) is computed; check and variable nodes such that $\nu(c_{4})\neq 0$ and $\gamma(c_{4})\neq 0$ are represented in gray, while nodes such that $\nu(c_{4})=0$ and $\gamma(c_{4})=0$ are represented in white.
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
(b) The influence of Fig. 1(a) is shifted to the node $c_{8}$: in this case, $\sigma_{x}=\sigma_{y}=1$.
Figure 2: Example of shifting of the proximity influence from $c_{5}$ to $c_{8}$.
IV-2 Auxiliary proximity influence
We also define an auxiliary proximity influence of a check node $c_{j}$, that we denote as $\mathbf{v}(c_{j})$.
Definition 3.
The auxiliary proximity influence $\mathbf{v}(c_{j})$ is a $1\times n$ vector such that $v_{i}(c_{j})$ is the length of the shortest path between the variable node $v_{i}$ and the check node $c_{j}$. For example, if a variable node $v_{i}$ is directly connected to $c_{j}$, then $v_{i}(c_{j})=1$.
Note that the proximity influences $\boldsymbol{\nu}(c_{j})$ and $\mathbf{v}(c_{j})$ share the same non-zero positions, but in general they have different values: while $\boldsymbol{\nu}(c_{j})$ will have higher values for the variable nodes close to $c_{j}$, in $\mathbf{v}(c_{j})$ the variable nodes connected to $c_{j}$ will have value $1$, those at distance 2 will have value $2$, and so on. We can compute $\mathbf{v}(c_{j})$ in a similar way to $\boldsymbol{\nu}(c_{j})$. Thus, it is sufficient to construct $\mathbf{v}(c_{1})$ offline, and apply Algorithm 1 to create $\mathbf{v}(c_{j})$, for any $j$. We illustrate in Section V-A1 how we make use of the auxiliary proximity influence to correct errors. Finally, we define two subroutines, illustrated in Algorithm 2 and Algorithm 3. The $\mathtt{Compute\_proximity\_vector}$ subroutine takes as input the syndrome $\mathbf{s}$ and generates the proximity vectors $\boldsymbol{\nu}$ and $\boldsymbol{\gamma}$, using Algorithm 1 to compute individual influences and summing them up according to (11). The $\mathtt{Shift\_and\_remove}$ subroutine takes as input the residual syndrome $\mathbf{s}^{\prime}$, as well as the proximity vectors $\boldsymbol{\nu}$ and $\boldsymbol{\gamma}$, and updates them using (12).
Algorithm 2 $\mathtt{Compute\_proximity\_vector}$
Input: $\mathbf{s}$
Global: $\boldsymbol{\gamma}(c_{1})$, $\boldsymbol{\nu}(c_{1})$, $L$
Output: $\boldsymbol{\nu}$, $\boldsymbol{\gamma}$
$\mathbf{k}\leftarrow k\ |\ s_{k}=1$
2:$\boldsymbol{\nu},\boldsymbol{\gamma}\leftarrow\mathbf{0}$
3:for $k\in\mathbf{k}$ do
4:     $\boldsymbol{\gamma}(c_{k}),\boldsymbol{\nu}(c_{k})\leftarrow\mathtt{Shift\_influence}(k)$
5:     $\boldsymbol{\gamma}\leftarrow\boldsymbol{\gamma}+\boldsymbol{\gamma}(c_{k})$
6:     $\boldsymbol{\nu}\leftarrow\boldsymbol{\nu}+\boldsymbol{\nu}(c_{k})$
7:end for
8:return
Algorithm 3 $\mathtt{Shift\_and\_remove}$
Input: $\boldsymbol{\nu}$, $\boldsymbol{\gamma}$, $\mathbf{s}^{\prime}$
Global: $\boldsymbol{\gamma}(c_{1})$, $\boldsymbol{\nu}(c_{1})$, $L$
Output: $\boldsymbol{\nu}^{\prime}$, $\boldsymbol{\gamma}^{\prime}$
$\mathbf{j}\leftarrow j\ |\ s^{\prime}_{j}=1$
2:$\boldsymbol{\nu}^{\prime}\leftarrow\boldsymbol{\nu}$
3:$\boldsymbol{\gamma}^{\prime}\leftarrow\boldsymbol{\gamma}$
4:for $k\in\mathbf{j}$ do
5:     $\boldsymbol{\gamma}(c_{k}),\boldsymbol{\nu}(c_{k})\leftarrow\mathtt{Shift\_influence}(k)$
6:     $\boldsymbol{\nu}^{\prime}\leftarrow\boldsymbol{\nu}^{\prime}-\boldsymbol{\nu}(c_{k})$
7:     $\boldsymbol{\gamma}^{\prime}\leftarrow\boldsymbol{\gamma}^{\prime}-\boldsymbol{\gamma}(c_{k})$
8:end for
9:return
V Progressive-Proximity Bit-Flipping
We are now ready to describe our proposed decoder, which we call Progressive-Proximity Bit-Flipping (PPBF). It is illustrated in Algorithm 4, and it is composed of two decoding steps: the first is called Preliminary BF, and the second is called Iterative matching; these two algorithms are illustrated in Algorithm 5 and Algorithm 6, respectively.
Algorithm 4 Progressive-Proximity Bit-Flipping
Input: $\mathbf{s}$, $\mathbf{H}$
Output: $\hat{\mathbf{e}}$
1:$\boldsymbol{\nu},\boldsymbol{\gamma}\leftarrow\mathtt{Compute\_proximity\_metric}(\mathbf{s})$
2:$\hat{\mathbf{e}},\hat{\mathbf{s}},\boldsymbol{\nu}^{\prime}\leftarrow\mathtt{Preliminary\_BF}(\mathbf{s},\mathbf{H},\boldsymbol{\nu})$
3:$\hat{\mathbf{e}}\leftarrow\mathtt{Iterative\_matching}(\hat{\mathbf{s}},\mathbf{H},\boldsymbol{\gamma})$
4:return
V-A Preliminary BF
In the preliminary BF step of Algorithm 5, the vector $\mathbf{u}$ containing the number of unsatisfied checks for each variable node is first computed (line 3), and only variable nodes involved in two unsatisfied checks will be considered for flipping. Among all the variable nodes $i$ such that $u_{i}=2$, we flip the one which achieves the minimum value of the proximity vector $\nu_{i}^{\prime}$ (lines 7-8); then, the proximity vector is updated using Algorithm 3, such that the influence of the checks satisfied after the flip is removed from $\boldsymbol{\nu}^{\prime}$ (line 10). The residual syndrome is then computed (line 11) and another iteration is performed, until there are no variable nodes $i$ such that $u_{i}=2$.
Algorithm 5 $\mathtt{Preliminary\_BF}$
Input: $\mathbf{s}$, $\mathbf{H}$, $\boldsymbol{\nu}$
Output: $\hat{\mathbf{e}}$, $\hat{\mathbf{s}}$, $\boldsymbol{\nu}^{\prime}$
1:$\hat{\mathbf{e}}\leftarrow\mathbf{0}$
[…]
3:$\mathbf{u}\leftarrow\mathbf{s}\cdot\mathbf{H}$
4:$\boldsymbol{\nu}^{\prime}\leftarrow\boldsymbol{\nu}$
5:$\mathcal{S}\leftarrow\{i\in[1,n]:u_{i}=2\}$
6:while $\mathcal{S}\neq\emptyset$ do
7:     $j\leftarrow\underset{i\in\mathcal{S}}{\mathrm{arg\ min}}\ \nu_{i}^{\prime}$
8:     $\hat{e}_{j}\leftarrow\hat{e}_{j}\oplus 1$
9:     $\mathbf{s}^{\prime}\leftarrow\hat{\mathbf{e}}\cdot\mathbf{H}^{T}\mod{2}$
10:     $\boldsymbol{\nu}^{\prime}\leftarrow\mathtt{Shift\_and\_remove(\boldsymbol{\nu}^{\prime},\cdot,\mathbf{s}^{\prime})}$
11:     $\hat{\mathbf{s}}\leftarrow\mathbf{s}\oplus\mathbf{s}^{\prime}$
12:     $\mathbf{u}\leftarrow\hat{\mathbf{s}}\cdot\mathbf{H}$
13:end while
14:return $\hat{\mathbf{e}}$, $\hat{\mathbf{s}}$
V-A1 The iterative-matching routine
[…]
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
(a) Proximity influence of depth 2 for the check node $c_{8}$; different coloring of variable nodes correspond to their distance from $c_{8}$: orange nodes are at distance 1, yellow at distance 2.
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
(b) Proximity influence of depth 2 for $c_{4}$.
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
(c) Combination of the previous two proximity influences: variable nodes which were assigned yellow and orange are now colored red; those nodes which were assigned yellow/yellow are now deep yellow. Finally, the variable nodes which were assigned orange/white or yellow/white have maintained their color. The set of red variable nodes coincide with the union of all the possible error patterns that satisfy the matching, i.e., $\{v_{10},v_{13}\}$ and $\{v_{7},v_{11}\}$.
Figure 3: Example of the usage of the auxiliary proximity influence of checks in Algorithm 6. In the example, $c_{4}$ and $c_{8}$ are unsatisfied checks, and the decoder has to find the shortest path between them, assuming to start with $c_{8}$ as pivot.
0000111144444444666633333333111122225555555577772222003333333388889999666612121212777710101010131313138888111111111414141466669999151515150000111116161616111122221717171722220015151515
Figure 4: Addition of the extra check as described in Section V-A1. The blue variable nodes are the depth-1 influence of the extra check, and the darker ones are those which will be flipped. In the example, $\Delta x=\Delta y=1$.
Algorithm 6 $\mathtt{Iterative\_matching}$
Input: $\hat{\mathbf{e}}$, $\mathbf{s}$, $\mathbf{H}$, $\boldsymbol{\gamma}$
Output: $\hat{\mathbf{e}}$
[…]
2:$\boldsymbol{\gamma}^{\prime}\leftarrow\boldsymbol{\gamma}$
3:while $|\hat{\mathbf{s}}|>0$ do
4:     $i\leftarrow\underset{i\in[1,m]}{\mathrm{arg\ min}}\ \boldsymbol{\gamma}^{\prime}\ |\ \hat{s}_{i}=1$triangleright Pivot.
5:     $\mathbf{t}\leftarrow\hat{\mathbf{s}}$
6:     $t_{i}\leftarrow 0$
7:     $\mathbf{c}(c_{i}),\mathbf{v}(c_{i})\leftarrow\mathtt{Shift\_influence}(i)$
8:     $\mathbf{j}\leftarrow j\ |\ t_{j}>0\wedge\mathbf{c}_{j}(c_{i})>0$
9:     $j\leftarrow\underset{j\in\mathbf{j}}{\mathrm{arg\ min}}\ \mathbf{c}(c_{i})\wedge\underset{j\in\mathbf{j}}{\mathrm{arg\ min}}\ \boldsymbol{\gamma}^{\prime}$triangleright Target.
10:     $\mathbf{c}(c_{j}),\mathbf{v}(c_{j})\leftarrow\mathtt{Shift\_influence}(j)$
11:     $\delta\leftarrow c_{j}(c_{i})$
12:     $\mathbf{f}\leftarrow k\ |\ v_{k}(c_{i})+v_{k}(c_{j})=\delta+1$
13:     if $|\mathbf{f}|=\delta$ then
14:         $\hat{e}_{\mathbf{f}}\leftarrow\hat{e}_{\mathbf{f}}\oplus 1$
15:     else
16:         if $j-i\mod L\ <i-j\mod L$ then
17:              $\Delta x\leftarrow j-i\mod L$
18:         else
19:              $\Delta x\leftarrow-(i-j\mod L)$
20:         end if
21:         $\Delta y\leftarrow\mathrm{min}\{\lfloor\frac{i-1}{L}\rfloor-\lfloor\frac{j-1}{L}\rfloor\mod L,\ \lfloor\frac{j-1}{L}\rfloor-\lfloor\frac{i-1}{L}\rfloor\mod L\}$
22:         $z\leftarrow i+\Delta x\mod L+L\lfloor\frac{i-1}{L}\rfloor$
23:         $\mathbf{c}(c_{z}),\mathbf{v}(c_{z})\leftarrow\mathtt{Shift\_influence}(z)$
24:         $\mathbf{f}_{1}\leftarrow k\ |\ v_{k}(c_{i})+v_{k}(c_{z})=\Delta x+1$
25:         $\mathbf{f}_{2}\leftarrow k\ |\ v_{k}(c_{j})+v_{k}(c_{z})=\Delta y+1$
26:         $\hat{e}_{\mathbf{f}_{1},\mathbf{f}_{2}}\leftarrow\hat{e}_{\mathbf{f}_{1},\mathbf{f}_{2}}\oplus 1$
27:     end if
28:     $\mathbf{s}^{\prime}\leftarrow\hat{\mathbf{e}}\cdot\mathbf{H}^{T}\mod{2}$
29:     $\boldsymbol{\gamma}^{\prime}\leftarrow\mathtt{Shift\_and\_remove}(\cdot,\boldsymbol{\gamma},\mathbf{s}^{\prime})$
30:     $\hat{\mathbf{s}}\leftarrow\mathbf{s}\oplus\mathbf{s}^{\prime}$
31:end while
32:return $\hat{\mathbf{e}}$
V-A2 Decoding of non-toric codes
It is possible to generalize the algorithm for non-toric surface codes, i.e., for those codes which present open borders, such that the variable nodes at the boundaries will have degree 1. An example that we present in this paper is the rotated planar code [28]; the X-Tanner graph of the $\llbracket 9,1,3\rrbracket$ rotated planar code is illustrated in Fig. 5.
In principle, our decoder is not suitable for non-toric codes, due to degree-1 variable nodes: indeed, if one of those variable nodes is in error, they can not be corrected by the BF because they only have degree 1, and neither by the iterative matching procedure, because there will only be one unsatisfied check, while we need at least two to perform a matching. Note that every matching decoder for a surface code needs to deal with this problem. To solve this issue, we adopt a similar solution solution to that proposed for MWPM in [2]: we attach virtual checks to the boundaries of the code, as shown in Fig. 5. The virtual checks are only considered in the iterative matching procedure; they are always considered to be unsatisfied, and the matching decoder treats them as normal checks, except for the fact that a virtual check cannot be chosen as a starting pivot.
00001111222211113333444455552222666677778888333344445555
Figure 5: Tanner graph for the $\llbracket 9,1,3\rrbracket$ rotated planar code. Black check nodes with dashed edges represent virtual check nodes.
Note that now, with open boundaries, we can still use Algorithms 1 and 2, but (14) has to be reformulated to take into account the open boundaries:
$\begin{cases}k^{\prime}_{c}=(k_{c}-\sigma_{x}+\lfloor(k_{c}-1)/L\rfloor-\sigma_{y}L)\ \mathrm{if}\ k^{\prime}_{c}\leq L^{2}\\ k^{\prime}_{v}=(k_{v}-\sigma_{x}+\lfloor(k_{v}-1)/L\rfloor-2\sigma_{y}L)\ \mathrm{if}\ k^{\prime}_{v}\leq 2L^{2}\\ k^{\prime}_{c},k^{\prime}_{v}=-1\ \mathrm{otherwise.}\end{cases}$ (15)
In this way, the proximity influence does not extend out of the graph. Also, for virtual check nodes we assign arbitrary proximity values; in other words, $\gamma_{k^{\prime}_{c}}^{\ell}(c_{j})=\gamma_{k_{c}}^{\ell}(c_{i})$ if $c_{i}$ is not an auxiliary check and $k_{c}^{\prime}\leq L^{2}$, $\gamma_{k^{\prime}_{c}}^{\ell}(c_{j})=0$ if $c_{i}$ is not an auxiliary check and $k_{c}^{\prime}=-1$, and $\gamma_{k^{\prime}_{c}}^{\ell}(c_{j})=C$, where $C$ is an arbitrary constant value, if $c_{i}$ is an auxiliary check. The rest of the algorithm remains the same.
VI Complexity analysis
In this section, we analyze the complexity of the proposed PPBF decoder. The computation of the proximity vector for $c_{1}$ is done offline and only once, thus it does not contribute to the computational complexity of the algorithm.
VI-A Complexity of Algorithms 1, 2 and 3
Algorithms 1, 2 and 3 are extensively used in every decoding iteration, therefore it is crucial that their computational complexity is low. Algorithm 1 simply applies (13) and (14) to shift the proximity influences; this can be done with complexity $\mathcal{O}(1)$. Algorithm 3 applies Algorithm 1 to update the proximity vector; since we update the proximity vector after every flip, Algorithm 1 is applied twice (every flip satisfies two check nodes) for every Algorithm 3 call. Thus, Algorithm 3 has the same order of complexity as Algorithm 1. Finally, since Algorithm 2 is applied just once offline to compute $\boldsymbol{\gamma}(c_{1})$ and $\boldsymbol{\nu}(c_{1})$, it does not contribute to the decoding complexity; in any case, similarly for Algorithm 3, it only involves calling Algorithm 1 $|\mathbf{s}|$ times, thus it has the same complexity. We then conclude that the computational complexity of Algorithms 1, 2 and 3 is $\mathcal{O}(1)$. We also want to stress that, since the entries of the proximity vectors are integers with bounded values, and since Algorithms 1, 2 and 3 only involve elementary operations such as integer summation, the memory usage and the runtime of the algorithm will generally be low compared to those of other algorithms with similar computational complexity, but that use floating point calculations.
VI-B BF decoding complexity
Assuming that comparing elements of $\mathbf{u}$ and $\boldsymbol{\nu}$ has a negligible impact on the complexity, the only meaningful contribution comes from Algorithm 3 and the computation of $\mathbf{u}$ (lines 10 and 12 of Algorithm 5), which both have complexity $\mathcal{O}(n)$.
VI-C Iterative matching complexity
In each iteration of Algorithm 6, all of the operations can be performed in $\mathcal{O}(1)$; since the procedure is repeated for each pair of unsatisfied checks, namely $|\mathbf{s}|/2$ times, making the complexity of the algorithm proportional to $\mathcal{O}(|\mathbf{s}|)$, and since $|\mathbf{s}|=\mathcal{O}(m)$ [22], $m$ being the number of check nodes, and since $m=\mathcal{O}(n)$ (for instance, for the toric code we have $m=n/2$) we have that the complexity of our decoder is $\mathcal{O}(n)$. The MWPM complexity is in the order of $\mathcal{O}(n^{12}\log n^{2})$, but efficient implementations are available that achieve very small runtime for relatively small values of $n$ [22]. The Union Find decoder [3] achieves an almost-linear complexity of $\mathcal{O}(\alpha(n)n)$, where $\alpha(n)$ is the inverse of Ackermann’s function [29], which has been shown to be linear for practical values of $n$.
VII Results
In this section we present simulation results for toric and rotated surface codes. We perform our simulation assuming a bit-flip channel and assume perfect syndrome measurements, and we compare our decoder with MWPM and UF. For our simulations we fix $D=\lfloor L/2\rfloor$. For each data point in the plotted curves, the simulation was run until either 100 logical errors were obtained or $10^{5}$ error vectors were processed. In Table I we compare the threshold of the proposed PPBF algorithm for the 2D toric code with those of MWPM and UF; the table also shows the computational complexity of each of these different decoders.
Decoder Threshold Complexity
MWPM $10.3\%$ [1] $\mathcal{O}(n^{3})$
UF $9.9\%$ [3] $\mathcal{O}(\alpha(n)n)$
PPBF (proposed) $7.5\%$ $\mathcal{O}(n)$
Table I: Threshold and complexity comparison among our decoder and other decoders present in the literature.
In Fig. 6 we plot simulation results of our decoder on toric codes over the bit-flip channel. As can be seen from the figure, the threshold is around $7.5\%$; to obtain the threshold for the depolarizing channel, it is sufficient to multiply the threshold value by 3/2 [17].
Figure 6: Performance of the proposed decoder on toric codes of different sizes, assuming a BSC channel. The threshold, which is around $7.5\%$, is highlighted.
In Fig. 7 we plot simulation results of our decoder on rotated planar codes over the BSC, using the technique described in Section V-A2. The threshold can be seen to occur around $7\%$.
Figure 7: Performance of our decoder on planar rotated codes of different sizes, assuming a BSC channel. We highlighted the threshold, which is around $7\%$.
In Fig. 8 we highlight the comparison between our decoder, traditional BF, MWPM and UF for the toric code with $L=13$. For the traditional BF, in each iteration we flip all of the bits involved in two unsatisfied checks, and we run it for a maximum of 100 iterations. As expected, MWPM, having higher complexity, achieves the best performance, both in terms of threshold and waterfall; the UF presents slightly worse performance than MWPM, although it has significantly lower complexity. Our decoder still achieves good performance, comparable to that of MWPM and UF, while achieving low complexity and latency; we also need to stress that our decoder is a hard-decision decoder, while MWPM and UF both utilize soft information. Comparing to the classical BF decoder, the PPBF shows a significantly lower error rate.
Figure 8: Performance comparison between our decoder, traditional BF, MWPM and UF for distance 13 toric code over BSC.
Finally, in Fig. 9 we plot the average number of decoding iterations for different crossover probabilities (denoted with $p$) as a function of the code’s blocklength of different toric codes; each data point is obtained by averaging the number of iterations used to decode $10^{5}$ error vectors sampled with fixed crossover probability. As shown in the previous section, we demonstrate that the number of decoding iterations grows linearly with the blocklength.
Figure 9: Average number of decoding iterations per number of physical qubits of the toric code for different crossover probabilities, denoted with $p$.
VIII Conclusion
We have presented a new decoder for toric and surface codes which is able to achieve a very good decoding performance with a low decoding complexity and latency. First, we have defined a novel heuristic called the proximity influence, which assigns weights to variable and check nodes in the neighborhood of an unsatisfied check; we also showed that the proximity influence of a check node $c_{j}$ can be efficiently obtained by appropriately permuting the influence of an arbitrary node $c_{1}$ which is computed offline and stored in memory. The total contribution of the influences of all of the unsatisfied checks is called the proximity vector, and we exploit it as a metric for flipping bits. To deal with error degeneracy and low variable node degrees of toric and surface codes, we designed the iterative matching procedure, which is employed after a round of serial BF. Future work could include improving the performance of the decoder using decoding diversity, i.e., running several rounds of PPBF, each one using a different variation of the proximity metric, and choosing as error estimate the one with least weight. Also, methods to incorporate soft information on the qubits or the syndrome into the proposed decoding algorithm could lead to performance improvement. Adapting the decoder to work on more general QLDPC codes could also could also be an interesting direction for future research.